# Element-Saving Hexahedral 3-Refinement Templates

### v1

@@ 2 Refinement Conditions
The greatest leap in subsequent research extended the capability of 3-refinement to handling concave edges by adopting a new set of face transition and cell transition templates [14]. Its face and cell transitions are shown in Figure 2 (B). Their key insight was that 1110 face could be locally decomposed into two 1100 face patterns. This decomposition enables transformation of 11100000 cell into two 11000000 cell templates via local refinement of its base face, and similarly facilitates conversion of 11101110 cell into two 11000000 cell templates and two 11110000 cell templates through local refinement of both base and top faces. However, this refinement is no longer recursive, as the quad faces attached to black dots are no longer unit squares. With only seven cell transition templates to cover the 22 possible vertex configurations, this method still relies on another refinement scheme in Table III of [14] to handle non-matching cases by adding black dots to map them onto the seven available templates. It can be readily verified that this template set is equivalent to enforcing the strongly-balanced condition while prohibiting concave corner transitions. However, in practical applications, refining cells intersected by the geometry surface inevitably introduces concave corner transitions within interior cavities. Consequently, like [25], this method forces uniform refinement of all interior cells to match the surface tree depth. The sole distinction is its ability to avoid subdividing concave edges in L-shaped sweeps, yielding only modest local cell count reduction without addressing the fundamental issue of over-refinement.
@@
Another method proposed later addresses vertex and edge irregularity while accommodating 3 concave transitions among the 22 possible vertex configurations [7]. As illustrated in Figure 2 (C), this method recursively assembles non-hex-meshable cells from meshable cells through localized refinement: 11100000 cell is decomposed into two 11000000 cells via corner refinement; 11100100 cell is decomposed into three 11100000 cells; 11111000 cell is decomposed into one 11110000 and two 11100000 cells; 11101110 cell is decomposed into two 11110000 cells via edge refinement; and 11111110 cell is decomposed into three 11110000 cells. However, these transitions exhibit inherent edge inconsistencies. For example, 01 edge in the 1100 face configuration appears as a single segment, whereas the same edge in the 1110 face is split into two segments. This inconsistency stems from local refinement isolation around non-black vertices or edges connecting two non-black vertices, occurring in five patterns: 11100000, 11100100, 11111000, 11101110, and 11111110. To enforce consistency, the authors developed an algorithm that first marks these non-black vertices and edges, then processes all local refinements in adjacent grid cells before applying cell templates. Since a grid cell may border with multiple non-black vertices and edges, necessitating repeated refinements that degrade mesh quality, the algorithm directly performs 1-to-27 refinement on cells adjacent to excessive numbers of non-black vertices or edges. Additionally, to constrain refinement patterns to the ten cell transition cases, the method requires the grid to be moderately-balanced and to satisfy the pairing condition, wherein subdividing a cell requires subdividing all 27 children of its parent. While this approach successfully resolves the concave corner limitations of [14] and reduces vertex/edge irregularities, its complicated refinement criteria demand numerous iterative cycles to converge, resulting in slow computation time. Furthermore, for meshes without extensive concave corner regions, these criteria often produce more hex elements than both [14] and [25].
@@ 3 Vertex-Based Templates
Considering the presence (1) or absence (0) of a black dot at each vertex of a cell, 22 distinct configurations are obtained after eliminating duplicates through rotational and mirror symmetries [21]. Figure 3 (A, B) presents the complete cell-level templates. Ten templates in Figure 3 (A) are fundamental cases derived from Figure 2 (B), each readily verified to comply with the face schemes in Figure 3 (C), where edges connecting vertices of types 00 and 01 remain unchanged while type-11 edges are trisected, also ensuring edge compatibility. Twelve templates in Figure 3 (B) are composite cases, handled by extending the local refinement approach of [14] from patterns 11100000 and 11101110 to all composite configurations. Each composite cell is subdivided into five or six sub-cells, whose vertex patterns either match a fundamental template or correspond to pattern 11100000. In the former case, the sub-cell is directly replaced by its matching fundamental template; in the latter, the 11100000 template is then used, and all its sub-cells align with fundamental templates. Since all resulting sub-cells conform to fundamental templates that satisfy the face transition schemes, the final hex mesh is guaranteed to be conforming. For instance, the 11111110 composite pattern is decomposed into three sub-cells of 11110000 pattern, two sub-cells of 11100000 pattern, and one sub-cell of 00000000 pattern. Each 11100000 sub-cell is subsequently refined into two 11000000 sub-cells, two 10000000 sub-cells, and two 00000000 sub-cells, all fundamental cases, thereby completing the template substitution.
For this method, each grid edge is either preserved or uniformly trisected, requiring only a moderately-balanced condition on the input grid. Furthermore, since the algorithm operates solely through cell substitution, its time complexity is $O(n)$ where $n$ denotes the number of grid cells, ensuring efficient and predictable performance. In summary, the proposed templates represent a substantial extension of [14] that significantly restricts refinement propagation. While certain cells may be subdivided into more elements locally, for instance, cell 11111110 yields 27 elements when refined to 11111111 using [14] but 55 elements with vertex-based templates, the constrained propagation ultimately produces fewer global hex elements. Numerical examples in Section 5 demonstrate that, despite occasional local increases, the vertex-based approach consistently generates fewer total elements than previous 3-refinement methods.
@@ 4 Comprehensive Edge-Based Templates
The template replacement method described in Section 3 adopts a vertex-based refinement criterion, consistent with prior approaches [25, 14, 30, 7]. This yields a compact set of only 22 templates, which facilitate clear presentation and straightforward implementation. However, the moderately-balanced condition requires only that edge transitions between neighboring cells be bounded by $1\rightarrow 3$ and face transitions by $1\times 1\rightarrow 3\times 3$. In essence, this permits each cell edge to comprise either one or three segments and each face to be either a single quad or a $3\times 3$ subdivision. Vertex-based templates, by contrast, impose stricter constraints than these conditions alone demand. As an example, consider the 1111 template face from Figure 3 (C): adjacency may occur only along the bottom and top edges with deeper-level cells. Yet under vertex-based marking, any vertex contacting a deeper-level cell must be marked as a black dot, thereby marking all four vertices and forcing the left and right edges into $1\rightarrow 3$ transitions, introducing unnecessary subdivision.
@@
As illustrated in the first 000000000000 cell in Figure 4 (A), a grid cell’s subdivision state can be encoded using twelve binary indicators $r(i)\in\{0,1\}$ for $i=1,2,\dots,12$, where “0” denotes an edge without trisection and “1” denotes an edge with trisection (bolded). The algorithm begins by examining simplified configurations leveraging symmetry. After exhaustive enumeration, the first 31 specialized cases (excluding the final [X][X][X] cell) are presented in Figure 4 (A), where “[1110]” represents a face with three 1-edges and one 0-edge, and “[X]” represents “XXXX”. The replacement procedure, as implemented in practice, proceeds hierarchically. The edge encoding of the cell under consideration is first matched against the first 29 templates among these 31 cases, specifically excluding the [1110][X][1110] and [1110][X][X] cells. Since the encodings for these 29 templates are uniquely determined, they require no prescribed processing order; upon locating a match in the edge encoding hash table, the corresponding replacement is applied immediately. If no match is found, the algorithm next evaluates the 30th template [1110][X][1110], which identifies configurations where a pair of opposite faces each contain three edges coded as “1” and one edge coded as “0”. This template isolates the four side faces into sub-cells, each of which is processed using the universal templates from Figure 4 (B) according to its side-face edge encoding. Subsequently, the 31st template [1110][X][X] is assessed, matching cells where a single face has three edges coded as “1” and one edge coded as “0”; it isolates the remaining five faces into sub-cells handled via Figure 4 (B). Finally, cells matching none of the first 31 templates are processed using the 32nd template [X][X][X], which isolates all six faces into sub-cells that are handled via Figure 4 (B). The reason why the last case can cover all the cells that do not match the first 31 templates is because in the most general case, such a cell can always be partitioned via local refinement into seven sub-cells. Up to rotational symmetry, the twelve-edge subdivision encodings for these sub-cells are: $0,0,0,0,\mathbf{0}$; $r(1),r(2),r(3),r(4),\mathbf{0}$; $r(11),r(10),r(9),r(12),\mathbf{0}$; $r(9),r(6),r(1),r(5),\mathbf{0}$; $r(3),r(7),r(11),r(8),\mathbf{0}$; $r(4),r(8),r(12),r(5),\mathbf{0}$; and $r(10),r(7),r(2),r(6),\mathbf{0}$, where $\mathbf{0}$” represents that the remaining eight bits are zero. Since the last eight bits of each encoding are invariably zero, the template design is simplified, requiring only the four edge states per face. This yields the self-consistent, edge-based face template scheme shown in Figure 4 (C). Applying these six face templates to a sub-cell in the last [X][X][X] cell of Figure 4 (C) produces six conforming templates in Figure 4 (B), thereby ensuring the effectiveness of the universal scheme. However, the universal scheme is inefficient, generating substantial redundant elements. For example for the 111111111111 cell, it produces one central 000000000000 cell and six face-corresponding 111100000000 cells, totaling 79 cells, far exceeding the optimal 27-cell subdivision. This inefficiency necessitates the design of the first 31 cell-efficient templates that exploit inherent symmetries for specialized cases. Consequently, conformal hex mesh templates can be constructed for all 4,096 possible edge subdivision configurations using this comprehensive approach.

### v2

@@ 2 Refinement Conditions
The greatest leap in subsequent research extended the capability of 3-refinement to handling concave edges by adopting a new set of face transition and cell transition templates [14]. Its face and cell transitions are shown in Figure 2 (B). Their key insight was that 1110 face could be locally decomposed into two 1100 face patterns. This decomposition enables transformation of 11100000 cell into two 11000000 cell templates via local refinement of its base face, and similarly facilitates conversion of $1110$ cell into two 11000000 cell templates and two 11110000 cell templates through local refinement of both base and top faces. However, this refinement is no longer recursive, as the quad faces attached to black dots are no longer unit squares. With only seven cell transition templates to cover the 22 possible vertex configurations, this method still relies on another refinement scheme in Table III of [14] to handle non-matching cases by adding black dots to map them onto the seven available templates. It can be readily verified that this template set is equivalent to enforcing the strongly-balanced condition while prohibiting concave corner transitions. However, in practical applications, refining cells intersected by the geometry surface inevitably introduces concave corner transitions within interior cavities. Consequently, like [25], this method forces uniform refinement of all interior cells to match the surface tree depth. The sole distinction is its ability to avoid subdividing concave edges in L-shaped sweeps, yielding only modest local cell count reduction without addressing the fundamental issue of over-refinement.
@@
Another method proposed later addresses vertex and edge irregularity while accommodating 3 concave transitions among the 22 possible vertex configurations [7]. As illustrated in Figure 2 (C), this method recursively assembles non-hex-meshable cells from meshable cells through localized refinement: 11100000 cell is decomposed into two 11000000 cells via corner refinement; 11100100 cell is decomposed into three 11100000 cells; 11111000 cell is decomposed into one 11110000 and two 11100000 cells; $1110$ cell is decomposed into two 11110000 cells via edge refinement; and 11111110 cell is decomposed into three 11110000 cells. However, these transitions exhibit inherent edge inconsistencies. For example, 01 edge in the 1100 face configuration appears as a single segment, whereas the same edge in the 1110 face is split into two segments. This inconsistency stems from local refinement isolation around non-black vertices or edges connecting two non-black vertices, occurring in five patterns: 11100000, 11100100, 11111000, $1110$, and 11111110. To enforce consistency, the authors developed an algorithm that first marks these non-black vertices and edges, then processes all local refinements in adjacent grid cells before applying cell templates. Since a grid cell may border with multiple non-black vertices and edges, necessitating repeated refinements that degrade mesh quality, the algorithm directly performs 1-to-27 refinement on cells adjacent to excessive numbers of non-black vertices or edges. Additionally, to constrain refinement patterns to the ten cell transition cases, the method requires the grid to be moderately-balanced and to satisfy the pairing condition, wherein subdividing a cell requires subdividing all 27 children of its parent. While this approach successfully resolves the concave corner limitations of [14] and reduces vertex/edge irregularities, its complicated refinement criteria demand numerous iterative cycles to converge, resulting in slow computation time. Furthermore, for meshes without extensive concave corner regions, these criteria often produce more hex elements than both [14] and [25].
@@ 3 Vertex-Based Templates
Considering the presence (1) or absence (0) of a black dot at each vertex of a cell, 22 distinct configurations are obtained after eliminating duplicates through rotational and mirror symmetries [21]. Figure 3 (A, B) presents the complete cell-level templates. Ten templates in Figure 3 (A) are fundamental cases derived from Figure 2 (B), each readily verified to comply with the face schemes in Figure 3 (C), where edges connecting vertices of types $0$ and 01 remain unchanged while type-11 edges are trisected, also ensuring edge compatibility. Twelve templates in Figure 3 (B) are composite cases, handled by extending the local refinement approach of [14] from patterns 11100000 and $1110$ to all composite configurations. Each composite cell is subdivided into five or six sub-cells, whose vertex patterns either match a fundamental template or correspond to pattern 11100000. In the former case, the sub-cell is directly replaced by its matching fundamental template; in the latter, the 11100000 template is then used, and all its sub-cells align with fundamental templates. Since all resulting sub-cells conform to fundamental templates that satisfy the face transition schemes, the final hex mesh is guaranteed to be conforming. For instance, the 11111110 composite pattern is decomposed into three sub-cells of 11110000 pattern, two sub-cells of 11100000 pattern, and one sub-cell of $0000$ pattern. Each 11100000 sub-cell is subsequently refined into two 11000000 sub-cells, two 10000000 sub-cells, and two $0000$ sub-cells, all fundamental cases, thereby completing the template substitution.
For this method, each grid edge is either preserved or uniformly trisected, requiring only a moderately-balanced condition on the input grid. Furthermore, since the algorithm operates solely through cell substitution, its time complexity is $O(n)$ where $n$ denotes the number of grid cells, ensuring efficient and predictable performance. In summary, the proposed templates represent a substantial extension of [14] that significantly restricts refinement propagation. While certain cells may be subdivided into more elements locally, for instance, cell 11111110 yields 27 elements when refined to $1111$ using [14] but 55 elements with vertex-based templates, the constrained propagation ultimately produces fewer global hex elements. Numerical examples in Section 5 demonstrate that, despite occasional local increases, the vertex-based approach consistently generates fewer total elements than previous 3-refinement methods.
@@ 4 Comprehensive Edge-Based Templates
The template replacement method described in Section 3 adopts a vertex-based refinement criterion, consistent with prior approaches [25, 14, 30, 7]. This yields a compact set of only 22 templates, which facilitate clear presentation and straightforward implementation. However, the moderately-balanced condition requires only that edge transitions between neighboring cells be bounded by $1\rightarrow 3$ and face transitions by $1\times 1\rightarrow 3\times 3$. In essence, this permits each cell edge to comprise either one or three segments and each face to be either a single quad or a $3\times 3$ subdivision. Vertex-based templates, by contrast, impose stricter constraints than these conditions alone demand. As an example, consider the $11$ template face from Figure 3 (C): adjacency may occur only along the bottom and top edges with deeper-level cells. Yet under vertex-based marking, any vertex contacting a deeper-level cell must be marked as a black dot, thereby marking all four vertices and forcing the left and right edges into $1\rightarrow 3$ transitions, introducing unnecessary subdivision.
@@
As illustrated in the first $000000$ cell in Figure 4 (A), a grid cell’s subdivision state can be encoded using twelve binary indicators $r(i)\in\{0,1\}$ for $i=1,2,\dots,12$, where “0” denotes an edge without trisection and “1” denotes an edge with trisection (bolded). The algorithm begins by examining simplified configurations leveraging symmetry. After exhaustive enumeration, the first 31 specialized cases (excluding the final [X]$[X]$ cell) are presented in Figure 4 (A), where “[1110]” represents a face with three 1-edges and one 0-edge, and “[X]” represents “XXXX”. The replacement procedure, as implemented in practice, proceeds hierarchically. The edge encoding of the cell under consideration is first matched against the first 29 templates among these 31 cases, specifically excluding the [1110][X][1110] and [1110]$[X]$ cells. Since the encodings for these 29 templates are uniquely determined, they require no prescribed processing order; upon locating a match in the edge encoding hash table, the corresponding replacement is applied immediately. If no match is found, the algorithm next evaluates the 30th template [1110][X][1110], which identifies configurations where a pair of opposite faces each contain three edges coded as “1” and one edge coded as “0”. This template isolates the four side faces into sub-cells, each of which is processed using the universal templates from Figure 4 (B) according to its side-face edge encoding. Subsequently, the 31st template [1110]$[X]$ is assessed, matching cells where a single face has three edges coded as “1” and one edge coded as “0”; it isolates the remaining five faces into sub-cells handled via Figure 4 (B). Finally, cells matching none of the first 31 templates are processed using the 32nd template [X]$[X]$, which isolates all six faces into sub-cells that are handled via Figure 4 (B). The reason why the last case can cover all the cells that do not match the first 31 templates is because in the most general case, such a cell can always be partitioned via local refinement into seven sub-cells. Up to rotational symmetry, the twelve-edge subdivision encodings for these sub-cells are: $0,0,0,0,\mathbf{0}$; $r(1),r(2),r(3),r(4),\mathbf{0}$; $r(11),r(10),r(9),r(12),\mathbf{0}$; $r(9),r(6),r(1),r(5),\mathbf{0}$; $r(3),r(7),r(11),r(8),\mathbf{0}$; $r(4),r(8),r(12),r(5),\mathbf{0}$; and $r(10),r(7),r(2),r(6),\mathbf{0}$, where $\mathbf{0}$” represents that the remaining eight bits are zero. Since the last eight bits of each encoding are invariably zero, the template design is simplified, requiring only the four edge states per face. This yields the self-consistent, edge-based face template scheme shown in Figure 4 (C). Applying these six face templates to a sub-cell in the last [X]$[X]$ cell of Figure 4 (C) produces six conforming templates in Figure 4 (B), thereby ensuring the effectiveness of the universal scheme. However, the universal scheme is inefficient, generating substantial redundant elements. For example for the $111111$ cell, it produces one central $000000$ cell and six face-corresponding 111100000000 cells, totaling 79 cells, far exceeding the optimal 27-cell subdivision. This inefficiency necessitates the design of the first 31 cell-efficient templates that exploit inherent symmetries for specialized cases. Consequently, conformal hex mesh templates can be constructed for all 4,096 possible edge subdivision configurations using this comprehensive approach.